$100 of savings 5%ps. over 3 years with a $1 monthly fee
1
Expert's answer
2017-04-25T13:26:08-0400
If we put $100 of savings with 5% per year over 3 years with a $1 monthly fee, then we will get
FV = PV*(1 + r)^n - fee*n*12 = 100*1.05^3 - 1*3*12 = $79.76.